# itemsep=1pt,topsep=2pt,parsep=1pt,partopsep=1pt

### v2

@@ 6.2. Fourth moment bounds for $\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi)$
The proof of Lemma 7.1 is similar to Theorem 6.3. We only need to obtain some moment bounds on $W_{s,t,u}^{\vec{\mathrm{i}},\vec{\mathrm{j}}}$. We postpone the proof of the lemma later. Note that in the proof of Theorem 6.3, we do not use the continuity of $\varphi$ and $\psi$ but only the boundedness of their values and supports. Hence, the above bound is also valid for $\varphi=\delta_{0}$. Also note that by assuming ${\mathbbm{E}}[\zeta_{n}^{3}]\geq 0$, the above fourth moment is increasing as the time interval getting larger. By (7.2), (7.14) and Lemma 7.1, we have that
$\begin{split}&\sum\limits_{\vec{\mathrm{i}}_{1}\in{\mathbbm{T}}_{\varepsilon},\vec{\mathrm{i}}_{2}\in A_{\vec{\mathrm{i}}_{1}},\vec{\mathrm{i}}_{3}\in A_{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}}}\frac{1}{3}\sum\limits_{k=1}^{3}\sup_{s,t,u}{\mathbbm{E}}\Big{[}\Big{|}\partial_{\vec{\mathrm{i}}_{k}}\mathscr{L}_{\varepsilon}\Big{(}W_{s,t,u}^{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}}\Big{)}\Big{|}^{4}\Big{]}^{\frac{3}{4}}\\ \leq&\sum\limits_{\vec{\mathrm{i}}_{1}\in{\mathbbm{T}}_{\varepsilon},\vec{\mathrm{i}}_{2}\in A_{\vec{\mathrm{i}}_{1}},\vec{\mathrm{i}}_{3}\in A_{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}}}C\bigg{\|}\frac{(\delta_{0})_{\varepsilon}}{w_{\varepsilon}}\bigg{\|}^{6}_{\ell^{p}}\|\psi\|_{\infty}^{6}\|w\mathbbm{1}_{B}\|^{6}_{\ell^{q}}\varepsilon^{\frac{3}{p}-\frac{3}{2}}\leq C_{\psi}\varepsilon^{\frac{3}{p}-\frac{5}{2}}(K_{\varepsilon})^{2}.\end{split}$
Hence, by choosing $1<p<\frac{6}{5}$, there exist some $C=C_{\psi,p}$ and $\delta>0$ such that $\limsup_{n\to\infty}|I_{1}|\leq C\varepsilon^{\delta}$, since $K_{\varepsilon}=(\log\frac{1}{\varepsilon})^{6}$, which concludes (7.5).
@@
(7.15) $\begin{split}|I_{1}^{(n)}|\leq&\frac{1}{2}\|f^{\prime\prime\prime}\|_{\infty}\sup_{\vec{\mathrm{i}}_{1}\in{\mathbbm{T}}_{\varepsilon}}{\mathbbm{E}}\big{[}|\Theta_{n}(\vec{\mathrm{i}}_{1})|^{3}\big{]}\sum\limits_{\vec{\mathrm{i}}_{1}\in{\mathbbm{T}}_{\varepsilon},\vec{\mathrm{i}}_{2}\in A_{\vec{\mathrm{i}}_{1}},\vec{\mathrm{i}}_{3}\in A_{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}}}\\ &\sup_{t,u}{\mathbbm{E}}\Big{[}\big{|}\partial_{\vec{\mathrm{i}}_{1}}\mathscr{L}_{\varepsilon}(W_{t,u}^{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}})\big{|}^{3}\Big{]}^{\frac{1}{3}}\sup_{t,u}{\mathbbm{E}}\Big{[}\big{|}\partial_{\vec{\mathrm{i}}_{2}}\mathscr{L}_{\varepsilon}(W_{t,u}^{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}})\big{|}^{3}\Big{]}^{\frac{1}{3}}\sup_{t,u}{\mathbbm{E}}\Big{[}\big{|}\partial_{\vec{\mathrm{i}}_{3}}\mathscr{L}_{\varepsilon}(W_{t,u}^{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}})\big{|}^{3}\Big{]}^{\frac{1}{3}},\end{split}$
@@
The proof are exactly the same as that for (7.5). We can find some constant $C\in(0,+\infty)$ and $\delta>0$, such that $\limsup_{n\to\infty}|I_{2}|\leq C\varepsilon^{\delta}$, which concludes the (7.6). Note that $W_{t,u}^{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}}$ and $W_{s,t,u}^{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}}$ have the same structure. In particular, both of them can be expressed by $c_{1}\Theta_{N}+c_{2}\Theta_{N}^{\rm(G)}$ with $c_{1},c_{2}\in[0,1]$ (see (7.11)). Hence, we can have a bound similar to Lemma 7.1, which is
@@
(7.16) $|I_{3}|\leq\frac{1}{2}\|f^{\prime\prime}\|_{\infty}\sum\limits_{\vec{\mathrm{i}}\in{\mathbbm{T}}_{\varepsilon}}\Big{(}{\mathbbm{E}}\big{[}\Theta_{n}^{2}(\vec{\mathrm{i}})\big{]}-{\mathbbm{E}}\big{[}\Theta_{m}^{2}(\vec{\mathrm{i}})\big{]}\Big{)}\sup\limits_{t\in[0,1]}{\mathbbm{E}}\big{[}|\partial_{\vec{\mathrm{i}}}\mathscr{L}_{\varepsilon}(W_{t})|^{2}\big{]},$
@@
First note that by Lemma 6.1, for any fixed $\varepsilon>0$, $\lim_{m,n\to\infty}\big{(}{\mathbbm{E}}[\Theta_{n}^{2}(\vec{\mathrm{i}})]-{\mathbbm{E}}[\Theta_{m}^{2}(\vec{\mathrm{i}})]\big{)}=0$. Then note that $\partial_{\vec{\mathrm{i}}}\mathscr{L}_{\varepsilon}(W_{t})$ is a multilinear polynomial in $W_{t}(\vec{\mathrm{i}})$ (see the discussion for (7.13)). Hence, ${\mathbbm{E}}[|\partial_{\vec{\mathrm{i}}}\mathscr{L}_{\varepsilon}(W_{t})|^{2}]$ is a multilinear polynomial in ${\mathbbm{E}}[W_{t}^{2}(\vec{\mathrm{i}})]$, which is a linear combination of ${\mathbbm{E}}[\Theta_{n}^{2}(\vec{\mathrm{i}})]$ and ${\mathbbm{E}}[\Theta_{m}^{2}(\vec{\mathrm{i}})]$ (we can assume that $\Theta_{n}^{\rm(G)}$ and $\Theta_{m}^{\rm(G)}$ are independent). Again by Lemma 6.1 and noting that there are finite many ${\mathbbm{E}}[W_{t}^{2}(\vec{\mathrm{i}})]$ for fixed $\varepsilon$, we have shown that $\lim_{n,m\to\infty}I_{3}^{(m,n)}=0$.
@@
We can actually extend Theorem 1.7 to a convergence in finite-dimensional distribution. For any fixed positive integer $k$, let $-\infty<s_{i}<t_{i}<+\infty$ and $\varphi_{i},\psi_{i}\in C_{c}({\mathbbm{R}})$ for $i=1,2,\cdots,k$, each component ${\mathcal{Z}}_{\lfloor s_{i}N\rfloor,\lfloor t_{i}N\rfloor}^{\beta_{N}}(\varphi_{i},\psi_{i})$ in the random vector $({\mathcal{Z}}_{\lfloor s_{i}N\rfloor,\lfloor t_{i}N\rfloor}^{\beta_{N}}(\varphi_{i},\psi_{i}))_{i=1,\cdots,k}$ can be approximated in $L_{2}$ by a coarse-grained model $\mathscr{L}_{\varepsilon,[s_{i},t_{i}]}(\varphi_{i},\psi_{i}|\Theta_{N,\varepsilon})$. We then have an $L_{2}$ approximation in vector level for fixed $k$. For any bounded test function $f:{\mathbbm{R}}^{k}\to{\mathbbm{R}}$ with bounded derivatives up to third order. The Lindeberg principle also holds for $f(\mathscr{L}_{\varepsilon,[s_{1},t_{1}]}(\cdot),\cdots,\mathscr{L}_{\varepsilon,[s_{k},t_{k}]}(\cdot))$ (see remark A.4). The proof for single ${\mathcal{Z}}_{N}$ can be applied here without any change.
@@
(7.17) $\begin{split}{\mathbf{M}}_{n,\varepsilon}^{\varphi,\psi}:&={\mathbbm{E}}\Big{[}\Big{(}\overline{\mathscr{L}}_{\varepsilon,[0,1]}\Big{(}\varphi,\psi|W_{s,t,u}^{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}\Big{)}\Big{)}^{4}\Big{]}=\varepsilon^{2}{\mathbbm{E}}\bigg{[}\bigg{(}\sum\limits_{r=1}^{(\log\frac{1}{\varepsilon})^{2}}\sum\limits_{(\vec{\mathrm{i}}_{1},\cdots,\vec{\mathrm{i}}_{r})\in\vec{\mathbf{{\mathcal{A}}}}_{\varepsilon}^{\text{\rm(no triple)}}}\sum\limits_{\mathrm{a},\mathrm{b}\in{\mathbbm{Z}}}\\ &\quad\quad\varphi_{\varepsilon}(\mathrm{a})g_{\mathrm{i}_{1}}(\mathrm{a})W_{s,t,u}^{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}(\vec{\mathrm{i}}_{1})\times\Big{\{}\prod_{j=2}^{r}g_{(\mathrm{i}_{j}-\mathrm{i}^{\prime}_{j-1})}(0)W_{s,t,u}^{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}(\vec{\mathrm{i}}_{j})\Big{\}}g_{(\lfloor\frac{1}{\varepsilon}\rfloor-\mathrm{i}^{\prime}_{r})}(\mathrm{b})\psi_{\varepsilon}(\mathrm{b})\bigg{)}^{4}\bigg{]}.\end{split}$
Recall the definition of $W_{s,t,u}^{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}$ from (7.11). We can write $W_{s,t,u}^{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}(\vec{\mathrm{i}})=c_{1}\Theta_{n}(\vec{\mathrm{i}})+c_{2}\Theta_{n}^{\rm(G)}(\vec{\mathrm{i}})$, where $c_{1}$ and $c_{2}$ are determined by i=($\vec{\mathrm{i}}=(\mathrm{i},\mathrm{i}^{\prime})$. Since $\Theta_{n}(\vec{\mathrm{i}})$ and $\Theta_{n}^{\rm(G)}(\vec{\mathrm{i}})$ are independent, the expectation of their products can be factorized. Similar to $\Theta_{n}(\vec{\mathrm{i}})$, $W_{s,t,u}^{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}(\vec{\mathrm{i}})$ and $W_{s,t,u}^{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}(\vec{\mathrm{j}})$ are uncorrelated for $\mathrm{i}\neq\mathrm{j}$. Also note that since the first and third moments of $\Theta_{n}^{\rm(G)}$ is 0, only its second and fourth moments
(7.18) ${\mathbbm{E}}\Big{[}\big{(}\Theta_{n}^{\rm(G)}(\vec{\mathrm{i}})\big{)}^{2}\Big{]}={\mathbbm{E}}\Big{[}\big{(}\Theta_{n}(\vec{\mathrm{i}})\big{)}^{2}\Big{]}\quad\text{and}\quad{\mathbbm{E}}\Big{[}\big{(}\Theta_{n}^{\rm(G)}(\vec{\mathrm{i}})\big{)}^{4}\Big{]}=3{\mathbbm{E}}\Big{[}\big{(}\Theta_{n}^{\rm(G)}(\vec{\mathrm{i}})\big{)}^{2}\Big{]}^{2}=3{\mathbbm{E}}\Big{[}\big{(}\Theta_{n}(\vec{\mathrm{i}})\big{)}^{2}\Big{]}^{2}$
can appear in the expansion. It means that to have a nonzero moment of some $\Theta_{n}^{G}(\vec{\mathrm{i}})$, $\vec{\mathrm{i}}$ must appear in the expansion for at least two times.
We can now adapt the same indices rearrangement in the proof of Theorem 6.3. Suppose that $1\leq n\leq\frac{2}{\varepsilon}$ is a renewal time. Then at least two mesoscopic renewal processes visit it, since this is necessary to have a nonzero moments of $\Theta_{n}$ or $\Theta_{n}^{\rm(G)}$. We can partition all renewal times into groups ${\mathcal{I}}_{1},\cdots,{\mathcal{I}}_{k}$. We are going to decouple $W_{s,t,u}^{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}(\vec{\mathrm{i}})$ with $\mathrm{i}\in{\mathcal{I}}_{s}$ and $\mathrm{i}^{\prime}\in{\mathcal{I}}_{t}$ for $t>s$.

### v3

@@ 6.1 Moment bounds for the coarse-grained variable $\Theta_{N,\varepsilon}$
The proof of Lemma 7.1 is similar to that of Theorem 6.3. We only need to obtain some moment bounds on $W_{s,t,u}^{\vec{\mathrm{i}},\vec{\mathrm{j}}}$. We postpone the proof of the lemma later. Note that in the proof for Theorem 6.3, we do not use the continuity of $\varphi$ and $\psi$ but only the boundedness of their values and supports (see Remark 4.2). Hence, the above bound is also valid for $\varphi(x)=\mathbbm{1}_{0}(x)$. Also note that by assuming ${\mathbbm{E}}[\zeta_{n}^{3}]\geq 0$, the above fourth moment is increasing as the time interval getting larger. By (7.2), (7.14) and Lemma 7.1, we have that
$\begin{split}&\sum\limits_{\vec{\mathrm{i}}_{1}\in{\mathbbm{T}}_{\varepsilon},\vec{\mathrm{i}}_{2}\in A_{\vec{\mathrm{i}}_{1}},\vec{\mathrm{i}}_{3}\in A_{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}}}\frac{1}{3}\sum\limits_{k=1}^{3}\sup_{s,t,u}{\mathbbm{E}}\Big[\Big|\partial_{\vec{\mathrm{i}}_{k}}\mathscr{L}_{\varepsilon}\Big(W_{s,t,u}^{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}}\Big)\Big|^{4}\Big]^{\frac{3}{4}}\\ \leq&\sum\limits_{\vec{\mathrm{i}}_{1}\in{\mathbbm{T}}_{\varepsilon},\vec{\mathrm{i}}_{2}\in A_{\vec{\mathrm{i}}_{1}},\vec{\mathrm{i}}_{3}\in A_{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}}}C\bigg\|\frac{(\delta_{0})_{\varepsilon}}{w_{\varepsilon}}\bigg\|^{6}_{\ell^{p}}\|\psi\|_{\infty}^{6}\|w\mathbbm{1}_{B}\|^{6}_{\ell^{q}}\varepsilon^{\frac{3}{p}-\frac{3}{2}}\leq C_{\psi}\varepsilon^{\frac{3}{p}-\frac{5}{2}}(K_{\varepsilon})^{2},\end{split}$
where $C_{\psi}$ does not depend on $B$, since $w(x)=e^{-|x|}$ (see Remark 6.4). Hence, by choosing $1<p<\frac{6}{5}$, there exist some $C=C_{\psi,p}$ and $\gamma>0$ such that $C\varepsilon^{\gamma}\geq\limsup_{n\to\infty}|I_{1}|$ by the choice of $K_{\varepsilon}$ in (2.35), which concludes (7.5).
@@
$\begin{split}|I_{1}^{(n)}|\leq&\frac{1}{2}\|f^{\prime\prime\prime}\|_{\infty}\sup_{\vec{\mathrm{i}}_{1}\in{\mathbbm{T}}_{\varepsilon}}{\mathbbm{E}}\big[|\Theta_{n}(\vec{\mathrm{i}}_{1})|^{3}\big]\sum\limits_{\vec{\mathrm{i}}_{1}\in{\mathbbm{T}}_{\varepsilon},\vec{\mathrm{i}}_{2}\in A_{\vec{\mathrm{i}}_{1}},\vec{\mathrm{i}}_{3}\in A_{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}}}\\ &\sup_{t,u}{\mathbbm{E}}\Big[\big|\partial_{\vec{\mathrm{i}}_{1}}\mathscr{L}_{\varepsilon}(W_{t,u}^{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}})\big|^{3}\Big]^{\frac{1}{3}}\sup_{t,u}{\mathbbm{E}}\Big[\big|\partial_{\vec{\mathrm{i}}_{2}}\mathscr{L}_{\varepsilon}(W_{t,u}^{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}})\big|^{3}\Big]^{\frac{1}{3}}\sup_{t,u}{\mathbbm{E}}\Big[\big|\partial_{\vec{\mathrm{i}}_{3}}\mathscr{L}_{\varepsilon}(W_{t,u}^{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}})\big|^{3}\Big]^{\frac{1}{3}},\end{split}$ (7.15)
@@
The proof is exactly the same as that for (7.5). We can find some constant $C\in(0,+\infty)$ and $\gamma>0$, such that $\limsup_{n\to\infty}|I_{2}|\leq C\varepsilon^{\gamma}$, which concludes (7.6). Note that $W_{t,u}^{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}}$ and $W_{s,t,u}^{\vec{\mathrm{i}}_{1},\vec{\mathrm{i}}_{2}}$ have the same structure. In particular, both of them can be expressed by $c_{1}\Theta_{N}+c_{2}\Theta_{N}^{\rm(G)}$ with $c_{1},c_{2}\in[0,1]$ (see (7.11)). Hence, we can have a bound similar to Lemma 7.1, which is
@@
$|I_{3}|\leq\frac{1}{2}\|f^{\prime\prime}\|_{\infty}\sum\limits_{\vec{\mathrm{i}}\in{\mathbbm{T}}_{\varepsilon}}\Big({\mathbbm{E}}\big[\Theta_{n}^{2}(\vec{\mathrm{i}})\big]-{\mathbbm{E}}\big[\Theta_{m}^{2}(\vec{\mathrm{i}})\big]\Big)\sup\limits_{t\in[0,1]}{\mathbbm{E}}\big[|\partial_{\vec{\mathrm{i}}}\mathscr{L}_{\varepsilon}(W_{t})|^{2}\big],$ (7.16)
@@
First note that by Lemma 6.1, for any fixed $\varepsilon>0$, $\lim_{m,n\to\infty}\big({\mathbbm{E}}[\Theta_{n}^{2}(\vec{\mathrm{i}})]-{\mathbbm{E}}[\Theta_{m}^{2}(\vec{\mathrm{i}})]\big)=0$. Then note that $\partial_{\vec{\mathrm{i}}}\mathscr{L}_{\varepsilon}(W_{t})$ is a multilinear polynomial of $W_{t}(\vec{\mathrm{i}})$ (see the discussion for (7.13)). Hence, ${\mathbbm{E}}[|\partial_{\vec{\mathrm{i}}}\mathscr{L}_{\varepsilon}(W_{t})|^{2}]$ is a multilinear polynomial of ${\mathbbm{E}}[W_{t}^{2}(\vec{\mathrm{i}})]$, which is a linear combination of ${\mathbbm{E}}[\Theta_{n}^{2}(\vec{\mathrm{i}})]$ and ${\mathbbm{E}}[\Theta_{m}^{2}(\vec{\mathrm{i}})]$ (we can assume that $\Theta_{n}^{\rm(G)}$ and $\Theta_{m}^{\rm(G)}$ are independent). Again by Lemma 6.1 and noting that there are finite many ${\mathbbm{E}}[W_{t}^{2}(\vec{\mathrm{i}})]$ for fixed $\varepsilon$, we have shown that $\lim_{n,m\to\infty}I_{3}^{(m,n)}=0$.
@@
We can actually extend Theorem 1.7 to a convergence in finite-dimensional distribution. For any fixed positive integer $k$, Let $\infty<s_{i}<t_{i}<+\infty$ and $\varphi_{i},\psi_{i}\in C_{c}({\mathbbm{R}})$ for $i=1,2,\cdots,k$. Each ${\mathcal{Z}}_{\lfloor s_{i}N\rfloor,\lfloor t_{i}N\rfloor}^{\beta_{N}}(\varphi_{i},\psi_{i})$ can be approximated in $L_{2}$ by a coarse-grained model $\mathscr{L}_{\varepsilon,[s_{i},t_{i}]}(\varphi_{i},\psi_{i}|\Theta_{N,\varepsilon})$. For the random vector $({\mathcal{Z}}_{\lfloor s_{i}N\rfloor,\lfloor t_{i}N\rfloor}^{\beta_{N}}(\varphi_{i},\psi_{i}))_{i=1,\cdots,k}$, we then have an $L_{2}$ approximation in vector level for fixed $k$. For any bounded test function $f:{\mathbbm{R}}^{k}\to{\mathbbm{R}}$ with bounded derivatives up to third order. The Lindeberg principle also holds for $f(\mathscr{L}_{\varepsilon,[s_{1},t_{1}]}(\cdot),\cdots,\mathscr{L}_{\varepsilon,[s_{k},t_{k}]}(\cdot))$ (see remark A.4). The proof for single ${\mathcal{Z}}_{N}$ can be applied here without any change.
@@
$\begin{split}{\mathbf{M}}_{n,\varepsilon}^{\varphi,\psi}:&={\mathbbm{E}}\Big[\Big(\overline{\mathscr{L}}_{\varepsilon,[0,1]}\Big(\varphi,\psi|W_{s,t,u}^{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}\Big)\Big)^{4}\Big]=\varepsilon^{2}{\mathbbm{E}}\bigg[\bigg(\sum\limits_{r=1}^{(\log\frac{1}{\varepsilon})^{2}}\sum\limits_{(\vec{\mathrm{i}}_{1},\cdots,\vec{\mathrm{i}}_{r})\in\vec{\mathbf{{\mathcal{A}}}}_{\varepsilon}^{\text{\rm(no triple)}}}\sum\limits_{\mathrm{a},\mathrm{b}\in{\mathbbm{Z}}}\\ &\quad\quad\varphi_{\varepsilon}(\mathrm{a})g_{\mathrm{i}_{1}}(\mathrm{a})W_{s,t,u}^{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}(\vec{\mathrm{i}}_{1})\times\Big\{\prod_{j=2}^{r}g_{(\mathrm{i}_{j}-\mathrm{i}^{\prime}_{j-1})}(0)W_{s,t,u}^{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}(\vec{\mathrm{i}}_{j})\Big\}g_{(\lfloor\frac{1}{\varepsilon}\rfloor-\mathrm{i}^{\prime}_{r})}(\mathrm{b})\psi_{\varepsilon}(\mathrm{b})\bigg)^{4}\bigg].\end{split}$ (7.17)
Recall the definition of $W_{s,t,u}^{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}$ from (7.11). We can write $W_{s,t,u}^{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}(\vec{\mathrm{i}})=c_{1}\Theta_{n}(\vec{\mathrm{i}})+c_{2}\Theta_{n}^{\rm(G)}(\vec{\mathrm{i}})$, where $c_{1}$ and $c_{2}$ are determined by $\vec{\mathrm{i}}=(\mathrm{i},\mathrm{i}^{\prime})$.
We can adapt the same indices rearrangement procedure in the proof of Theorem 6.3. Since the centered Gaussian family $\Theta^{\rm(G)}$ has the same covariance structure as $\Theta$, and since $\Theta^{\rm(G)}$ and $\Theta$ are independent, we have that for $\mathrm{i}\neq\mathrm{j}$, ${\mathbbm{E}}[W_{s,t,u}^{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}(\vec{\mathrm{i}})W_{s,t,u}^{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}(\vec{\mathrm{j}})]=0$. Hence, the argument in “summation rearrangement” part in the proof of Theorem 6.3 can be transferred here without change.
Then we need to check that the decoupling procedure, that is, the step (A) in “adjustments on coarse-grained disorders and heat kernels” part in the proof of Theorem 6.3 is still valid. We need to check that for $|\vec{\mathrm{i}}|\geq 2$, the second to fourth moments of $W_{s,t,u}^{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}(\vec{\mathrm{i}})$ can be controlled by the corresponding moments of $W_{s,t,u}^{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}((\mathrm{i},\mathrm{i}))g_{\mathrm{i}^{\prime}-\mathrm{i}}(0)W_{s,t,u}^{\vec{\mathrm{j}}_{1},\vec{\mathrm{j}}_{2}}((\mathrm{i}^{\prime},\mathrm{i}^{\prime}))$.